# Mitigate Position Bias in Large Language Models via Scaling a Single Dimension

### v2

@@ A.1 Datasets Details
We choose NaturalQuestion Multi-document QA and Key-Value Retrieval datasets used in "lost in the middle" paper (Liu et al., 2024c) to evaluate the degree to which our method alleviates position bias. NaturalQuestion Multi-document QA require the model to answer the question based on one key information document which is inserted in a long context consisting of many irrelevant documents. And Key-Value Retrieval needs the model to retrieve the value corresponding to the given key from a list consisting of hundreds of Key-Value pairs. These two datasets are both classic in-context tasks which aim to evaluate the differences of model performance when key information is located at different positions in the context. The evaluation metric is accuracy, based on whether the model’s response contains a string of the correct answer. In addition, we evaluate our method’s improvements across multi task types, using LongBench (Bai et al., 2024), a benchmark for bilingual, multitask, and comprehensive assessment of long context understanding capabilities of LLMs. It contains six major categories, covering single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks and code completion. The evaluation metrics are: F1 for single-document QA and multi-document QA, Rouge-L for summarization, accuracy (exact match) for few-shot learning and synthetic tasks, and edit similarity for code completion. During inference, since the original context may sometimes be too long, the input sequences will be truncated in the middle part to avoid exceeding the context window of the model.
A.2 Additional Implemention Details
Curve Fitting
@@ A.2 Additional Implemention Details
When we perform curve fitting on $h(p)$, we use least-squares cubic polynomial fit. And when judging its monotonicity, we skip the first 100 positions because the first a few values are often outliers. Since $h(p)$ is originally a discrete function, in practice, we employ the second-order difference to approximate the second-order derivative when computing smoothness.
Ms-PoE on Mistral
@@
Ablation of the Searching Algorithm
We conducted ablation experiments to demonstrate the necessity of using the three indicators (monotonicity, smoothness, validation loss) in our searching algorithm. Ours w/o monotonicity means we just select top-10 smoothest dimensions and then use the validation loss to determine. Ours w/o smoothness means we select top-10 dimensions with the highest number of monotonic layers and then use validation loss. Ours w/o validation loss means we first select top-10 dimensions with the highest number of monotonic layers and then just choose the smoothest one among them.
A.3 Scaled Dimensions Details
Table 5: The scaled dimensions, scale factors and applied layers of models.
Model Dimension Scale factor Applied layers
@@ A.3 Scaled Dimensions Details
The scaled dimensions, scale factors and applied layers of each model we use in out experiments are shown in Table 5.
A.4 Inference Latency
Table 6: Time consumed (minutes) of LLaMA-2-7b-chat in a single A100.
Method KV Retrieval NaturalQuestion
@@ A.4 Inference Latency
Appendix B Obtain Attention to Key Information
@@
The last token of the prompt will directly take on the task of predicting the answer, i.e., the value which need to be retrieved. Hence, the last token’s attention weights to the previous text can reflect whether it accurately retrieves the key information. We define the model’s attention (in some layer) to the key information as $A_{G}$ in Eq 6, where $G$ represents the set of token positions corresponding to where the key information is at, $l$ is the position of the last token of the prompt, and $a_{l,j}$ represents the attention weight of the $l$-th token to the $j$-th token. By shifting $G$, we use the same method to calculate its attention to each other KV pairs.
@@
Appendix C How We Modify Causal Mask and Position Embedding in KV Retrieval
In the method 1 in section 2.2, we crop the causal mask to let the "key tokens" unable to attend the previous tokens. As shown in Figure 7, the white part represents the cropped part, which means attention weights are 0, and the orange part represents the attention between tokens within key tokens. In addition, we have retained the attention of key tokens to the first token to maintain the stability of attention distribution. What is more, we only modify the causal mask in layers 1~8, but as the results, the attention to the key information is still significantly improved in layers 15~31, which indicates the positional information generated by causal mask in former layers can be transmitted to latter layers using posisional hidden states as the medium, thus modifying the causal mask solely in the former layers can induce a profound shift in the model’s comprehension of positional information.
@@ Appendix C How We Modify Causal Mask and Position Embedding in KV Retrieval
Appendix D Perturbation on Causal Mask and Position Embedding
@@
(c) Reduce PE
Figure 9: We performed perturbation experiments on the causal mask and position embedding (PE), showing the dimension 213 of hidden states of Mistral-7b (Jiang et al., 2023) using randomly synthesized corpus as input.
@@
Appendix E Attention Distribution Layer-wise and Head-wise
Figure 10 shows Mistral-7b’s attention to each KV pair of each layer (average across all attention heads) in the context in a KV retrieval task when the gold KV is put at different positions. The y-axis is the gold KV’s position, x-axis is each KV’s position, and the scale of the colorbar represents attention ($10^{-3}$). We can observe that diagonal patterns, which indicates the attention is concentrated on the "key tokens", appear only in the latter layers (start from layer 14), and may be a manifestation of retrieval behavior. In contrast, the former layers only focus on the beginning or end, regardless of where the key information is located.
Figure 11 shows the head-wise situation of layer 15. We can see actually only a portion of attention heads exhibit diagonal patterns, which may correspond to retrieval heads (Wu et al., 2024). The attention distribution in these heads also shows a pattern corresponding "loss in the middle", being larger at the beginning or end while significantly smaller at the middle.
@@ Appendix E Attention Distribution Layer-wise and Head-wise
Appendix F Positional Hidden States Visualization
(a) LLaMA-2-7b-chat dim=2393
(b) mistral-7b dim=213
(c) MPT-30b dim=6926
(d) Tinyllama-NoPE dim=1156
Figure 12: Positional hidden states output by each layer of LLaMA-2-7b-chat, Mistral-7b-Instruct-v0.2, MPT-30b-chat and TinyLlama-NoPE-1.1B. The x-axis represents the position, and the y-axis represents the value of the states.
We shown various models’ positional hidden states of each layer in Figure 12. When visualizing, we discarded the first 30 tokens because the hidden states values of these tokens are often huge (usually hundreds of times larger than the normal value (Sun et al., 2024)), which can disrupt monotonicity. We observed its monotonic trend first appears just in the first layer (actually just after the first attention mechanism), and continue to be more marked.

### v3

@@ Appendix B Datasets Details
We choose NaturalQuestion Multi-document QA and Key-Value Retrieval datasets used in “lost in the middle” paper (Liu et al., 2024) to evaluate the degree to which our method alleviates position bias. NaturalQuestion Multi-document QA requires the model to answer the question based on one key information document which is inserted in a long context consisting of many irrelevant documents. And Key-Value Retrieval needs the model to retrieve the value corresponding to the given key from a list consisting of hundreds of Key-Value pairs. These two datasets are both classic long-context tasks which aim to evaluate the differences of model performance when key information is located at different positions in the context. The evaluation metric is accuracy, based on whether the model’s response contains the string of the correct answer.
In addition, we evaluate our method’s improvements across multi task types, using LongBench (Bai et al., 2024), a benchmark for bilingual, multitask, and comprehensive assessment of long context understanding capabilities of LLMs. It contains six major categories, covering single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks and code completion. The evaluation metrics are: F1 for single-document QA and multi-document QA, Rouge-L for summarization, accuracy (exact match) for few-shot learning and synthetic tasks, and edit similarity for code completion. During inference, since the original context may sometimes be too long, the input sequences will be truncated in the middle part to avoid exceeding the context window of the model.
Appendix C Other Implementation Details
C.1 Curve Fitting
@@
When we perform curve fitting on $h(p)$, we use least-squares cubic polynomial fit. And when judging its monotonicity, we skip the first 30 positions because the first a few values are often outliers. Since $h(p)$ is originally a discrete function, in practice, we employ the second-order difference to approximate the second-order derivative when computing smoothness.
C.2 Ms-PoE on Mistral
@@
C.3 Parameters of the Scaling Method
Model Channel Index Scale factor Applied layers
@@ C.3 Parameters of the Scaling Method
Table 4: The scaled channels, scale factors and applied layers of models.
The scaled channel indices, scale factors and applied layers of each model we use in out experiments are shown in Table 4.
C.4 Inference Latency
Method KV Retrieval Multi-Doc
@@ C.4 Inference Latency
Table 5: Average inference time (minutes) of LLaMA-2-7b-chat in a single A100 on KV retrieval.
@@
Appendix D Details of The Confirmatory Experiments
D.1 Obtain Attention to Key Information
@@
The last token of the prompt will directly account for predicting the answer, i.e., the value which need to be retrieved. Hence, the last token’s attention weights to the previous text can reflect whether it accurately retrieves the key information. We define the model’s attention (in some layer) to the key information as $A_{G}$ in Eq 4, where $G$ represents the set of token positions corresponding to where the key information is at, $l$ is the position of the last token of the prompt, and $a_{l,j}$ represents the attention weight of the $l$-th token to the $j$-th token. By shifting $G$, we use the same method to calculate its attention to each other KV pairs.
@@
D.2 Attention is Related to Performance
(a) Attention
(b) Attention w/ modification
(c) Accuracy
(d) Accuracy w/ modification
Figure 7: Distribution of attention weight and accuracy as the gold KV is placed at different positions in the prompt. (b) and (d) are situations when the attention on the 25th KV pair is manually modified.
As shown in Figure 7, when we manually double the attention weights of the gold KV (i.e. the 25th KV, as illustrated in Figure 7b) during the model’s forward pass on the KV retrieval task, the retrieval accuracy when the gold KV is the 25th KV improves, (shown in Figure 7d). This demonstrates that the attention weights to the key information are positively correlated with retrieval accuracy.
D.3 How We Modify Causal Mask and Position Embedding in KV Retrieval
In the method 1 in section 2.2, we crop the causal mask to let the “key tokens” unable to attend the previous tokens. As shown in Figure 8, the white part represents the cropped part, which means attention weights are 0, and the orange part represents the attention between tokens within key tokens. In addition, we have retained the attention of key tokens to the first token to maintain the stability of attention distribution. What is more, we only modify the causal mask in layers 1~8, but as the results, the attention to the key information is still significantly improved in layers 15~31, which indicates the positional information generated by causal mask in former layers can be transmitted to latter layers using posisional hidden states as the medium, thus modifying the causal mask solely in the former layers can induce a profound shift in the model’s comprehension of positional information.
@@ D.3 How We Modify Causal Mask and Position Embedding in KV Retrieval
(a) Hidden States
(b) Attention
(c) Hidden States
(d) Attention
Figure 10: The positional hidden states and attention of each KV in a KV retrieval task (the gold KV is the 26th KV) of Mistral-7b-v0.2. (a)(b) We modify the hidden states of the 26th KV. (c)(d) We modify the hidden states of the 36th KV. (a)(c) The value of the 213rd channel of the hidden states (averaged across layer 15 to 20) of each KV (averaged across the tokens of the KV). (b)(d) The attention (averaged across layer 15 to 20 and across every attention head) to each KV.
D.4 Perturbation on Causal Mask and Position Embedding
@@
(c) Subtract position ids of PE
Figure 11: We performed perturbation experiments on the causal mask and position embedding (PE), showing the 213rd channel of hidden states of some layers of Mistral-7b (Jiang et al., 2023) using randomly synthesized corpus as input. It is clear that only cropping mask cause significant drop of positional hidden states values.
@@
D.5 Positional Hidden States Affect Attention
Although it is intuitive that these monotonically changing hidden states channels convey rich positional information, it is unclear how much impact one position channel will have on position bias. To find out this, we modify the position hidden states values to see corresponding attention changes, in a KV retrieval task with 50 KV pairs while the gold KV is the 26th KV.
Specifically, we subtract 0.3 from the 213rd channel of the hidden states of the tokens of the 26th and 36th KV respectively (in layer 15 to 20 of Mistral-7b-v0.2), and observe the change of positional hidden states (the 213rd channel) and attention weights of each KV.
As shown in Figure 10, we show each KV’s average hidden state value of the 213rd channel of Mistral-7b-v0.2 and the average attention to each KV, which are both averaged across layer 15 to 20 (because they are retrieval-related layers). It is clear that when the positional hidden states values of the 26th or 36th KV greatly drop, the attention to it is greatly increased, even over the level of the 1st KV, regardless of whether it is the gold KV. Therefore, we confirm only changing one positional channel can significantly affect attention, thus affect position bias.
Appendix E Ablation Experiments of the Searching Algorithm
Method LLaMA-2-7b Vicuna-13b Gemma-7b Mistral-7b Qwn1.5-7b
not applied 31.3 50.2 28.0 59.5 56.8
Ours 40.6 52.7 34.6 60.9 58.3
w/o monotonicity 40.6 51.8 34.6 60.9 58.3
w/o smoothness 40.6 52.7 27.8 60.9 58.3
w/o validation set 30.1 51.8 26.5 60.9 58.3
w/ scale top-2 channels 37.2 50.8 31.7 60.1 57.2
w/ modify last 16 tokens 41.6 51.5 34.6 59.7 58.1
w/ modify all tokens 44.0 50.8 31.7 59.5 57.4
Table 6: Average performance of different ground-truth positions using different methods on NaturalQuestions multi-document QA dataset (20 documents) Liu et al. (2024).
To evaluate the contributions of different components in our method, we introduce the following sets for the ablation study: (1) Ours w/o monotonicity, w/o smoothness, and w/o validation set, which adjust the search algorithm by not considering one of these three indicators, respectively (details in Appendix C). (2) Ours w/ scale 2 channels, which scales the top-2 positional hidden states simultaneously. (3) Ours w/ modify last 16 tokens and w/ modify all tokens, which adjust the range of tokens affected by the scaling operation in Equ.(3).
Table 6 shows the ablation results. It can be seen that without filtering by monotonicity or smoothness, performance may decline, and removing the validation set results in more decline in model performance. When the range of tokens or channels affected by scaling is expanded, most models experience varying degrees of performance loss. Considering these factors, we choose to modify only the last token and the top-1 positional channel to achieve the best performance.
Appendix F Attention Distribution Layer-wise and Head-wise
Figure 12 shows Mistral-7b’s attention to each KV pair of each layer (average across all attention heads) in the context in a KV retrieval task when the gold KV is put at different positions. The y-axis is the gold KV’s position, x-axis is each KV’s position, and the scale of the colorbar represents attention ($10^{-3}$). We can observe that diagonal patterns, which indicates the attention is concentrated on the “key tokens”, appear only in the latter layers (start from layer 14), and may be a manifestation of retrieval behavior. In contrast, the former layers only focus on the beginning or end, regardless of where the key information is located.
Figure 13 shows the head-wise situation of layer 15. We can see actually only a portion of attention heads exhibit diagonal patterns, which may correspond to retrieval heads (Wu et al., 2025). The attention distribution in these heads also shows a pattern corresponding “lost in the middle”, being larger at the beginning or end while significantly smaller at the middle.
@@ Appendix F Attention Distribution Layer-wise and Head-wise
Appendix G Positional Hidden States Visualization
(a) mistral-7b
(b) LLaMA-2-7b
(c) MPT-30b
(d) Tinyllama-NoPE
Figure 14: Hidden states values with the token positions of the positional channel of each layer. The x-axis represents the position, and the y-axis represents the value of the states. (a) The 213rd channel of Mistral-7b-v0.2 (b) The 2393th channel of Llama2-7b (c) The 1942th channel of MPT-30b (d) The 1156th channel of TinyLlama-NoPE-1.1B, which is a model without any position embeddings.
We show various models’ positional hidden states of each layer in Figure 14. When visualizing, we discard the first 30 tokens because the hidden states values of these tokens are often too huge (usually hundreds of times larger than the normal value (Sun et al., 2024)), which can disrupt monotonicity. We observe its monotonic trend may first appears just in the first layer (actually just after the first attention module), and continues to be more marked.